OpenStudy (anonymous):

Evaluate and write your answer in scientific notation: (4x10^2)(10*10^3) @-------College Al…

OpenStudy (anonymous):

\[4\times 3=12\] \[10^3\times 10^2=10^5\] so initially you get \[12\times 10^5\] in scientific notation this is \[1.2\times 10^6\]